OpenStudy (anonymous):

Prove that each equation is not a trigonometric identity by producing a counterexample. I will type

OpenStudy (anonymous):

\[\sec ^{2}x-1=\frac{\cos x}{\csc x}\]

OpenStudy (jamesj):

Try a few different values for x and find one which does not satisfy the equation; e.g., \[x = \pi/2, \ \pi/3, \ \pi/6, \ \pi/4\] At least one of these does not satisfy the equation.